the summation of ages of father and son is 60 years. what will the summation of their ages after 15 years?

Answer by josgarithmetic(39620)

f father age
s son age

f+s=60.
15 years in the future:
(f+15)+(s+15)= ?
f+15+s+15
f+s+30
(f+s)+30
60+30=90